OpenStudy (akashdeepdeb):

HELP NEEDED: If X = {8^n - 7n - 1|n ∈ N} and Y = {49(n-1)|n ∈ N}, then prove that X ⊆ Y. Do we just have to Plug-in values for n or is there any alternative method?

OpenStudy (akashdeepdeb):

@satellite73 @experimentX

OpenStudy (akashdeepdeb):

Anyone? :|

OpenStudy (anonymous):

lets see what we can do somehow it must be true that if \(z=8^n - 7n - 1\) then \(z\) can be written as \(49(n-1)\)

OpenStudy (akashdeepdeb):

Yeah

OpenStudy (akashdeepdeb):

Can we do this By combination?

OpenStudy (anonymous):

i don't know i am not that fast my first thought was to factor \(8^n-1\) as the difference of two cubes

OpenStudy (anonymous):

but now i think @experimentX has a better idea

OpenStudy (experimentx):

lol ... no i don't I just came back :) after walk

OpenStudy (akashdeepdeb):

heheh

OpenStudy (akashdeepdeb):

Thanks @satellite73 ! I'll go through it once more and see!

OpenStudy (anonymous):

maybe it would be easier by induction?

OpenStudy (akashdeepdeb):

Principal Of Mathematical Induction? I'll try THAT out! Thanks!! :D

OpenStudy (anonymous):

i will bet induction and some algebra will work certainly true for \(n=1\)

OpenStudy (experimentx):

show that \[ 1 + 8 + ... +8^{n-1} = n \mod 7 \]

OpenStudy (akashdeepdeb):

It DOES work with 1 and 2

OpenStudy (experimentx):

for all n , 8^n = 1 mod 7 add it n times from 0 to n-1

OpenStudy (anonymous):

Certainly works by induction: you can prove that \(8^n - 7n - 1\) is divisible by 49 for any \(n\). But experimentX's method looks neater...

OpenStudy (anonymous):

k but how to prove that \(8^k\equiv 1 (7)\)?

OpenStudy (anonymous):

and you are not allowed to say "by induction!"

OpenStudy (experimentx):

ab mod c = ((a mod c)(b mod c)) mod c

OpenStudy (anonymous):

yeah, but what about the \(k\)? ok i am being silly

OpenStudy (experimentx):

hmm ... (abcd ... k) mod c = (a mod c)(b mod c) ... (k mod c) mod c since all terms are 8 and the remainder is 1, the remainder is also 1. I think this can be done better via euler totient function. I am little out of touch on this.

OpenStudy (anonymous):

@satellite73 I guess your point is that 8^n = 1 mod 7 implicitly requires induction (e.g. ab mod c = ((a mod c)(b mod c)) mod c applied "repeatedly") to be proved formally? ;-)

OpenStudy (akashdeepdeb):

Thanks for your help. I got it by using Principal of Mathematical Induction.